Report of Activities on the RVIB N.B. Palmer Cruise 02-04  

21 August 2002


The survey work in the central sector of the Southern Ocean GLOBEC grid continued on 21 August with the vessel finishing work at station 44, steaming to station 45, and starting work there.  Both of these stations are furthest from the shore with station 44 at the end of survey line 6 and station 45 at the end of survey line 7.  Both were in the northeasterly flow of the Antarctic Circumpolar current. 


A tow with the 10-m MOCNESS went well except for a couple of snags of the towing wire on ice, but the catches were good and Jose Torres was especially pleased to have caught a fish species that he had never seen before and a couple of other rare ones.  This tow ended around 0330 and a little after 0400 the ice collectors were out on a floe next to the ship.  The ROV went in for a few minutes, but was sidelined by thruster problems. A 2000 m CTD cast was completed around 0800.  The 1-m MOCNESS tow, which was last up at station 44, was aborted because of a problem with the A-frame.  It was originally thought to be due to frozen hydraulic lines, but it turned out to be a pump motor that had failed and had to be replaced.  The replacement took several hours, delaying the deployment of BIOMAPER-II for the steam to station 45. Around 1330, the towed body was put into the water and was towyoed the remaining distance to the station.  It was an excellent day for seabird and marine mammal observations which took place throughout much of the daylight period.


The work at station 45 began at 1830 with the ice collectors being deployed onto a floe next to the ship, but the ROV was still under repair and was not deployed at the same time as had become the custom. This was followed by another 2000 m CTD cast and a Tucker trawl to collect live animals. Furcilia of Euphausia superba are in short supply this year and most stations now have Tucker trawls scheduled to try to collect enough for the shipboard experimental work.


The deep freeze that began on 19 August continued into a third day with temperature hovering around -25ºC.  The barometric pressure settled in around 984 mb and the winds were out of the southwest about 15 to 20 kts throughout the day.  About the time the CTD cast was being completed, the full moon was setting and the sun was rising.  Its light was filtered into a wondrous array of colors and reflections by sea smoke coming up from the open leads spread throughout the pack ice.  The exceptionally clear skies stayed the day. At the close of day, there were some wispy clouds near the horizon, which together with the sea smoke gave rise to the great sunset.  The pack ice was mostly 10/10 and there were a number of frozen leads - large flat newly formed ice areas with no snow on them.


CTD Group report (Eileen Hofmann, Bob Beardsley, Baris Salihoglu, Chris MacKay, Francisco (Chico) Viddi, Sue Beardsley)

During 21 August, we completed two CTD casts at stations 44 and 45, which are at the outer end of survey transect 7.  These stations were nominally in 865 m and 2966 m of water, respectively.  The actual location of station 44 is on the outer shelf slope, but an extended MOC-10 tow prior to the CTD cast resulted in the station being moved off the shelf.  As a result, this station ended up in 3410 m of water. Station 45, which is intended to be in deep water, was 2971 m deep. The consequence of moving station 44 off the shelf is that we obtained two CTD casts in oceanic water on the offshore side of the southern boundary of the Antarctic Circumpolar Current.


The CMiPS instrument remained on the CTD for both casts, which meant that the casts only went to 2000 m rather than the full depth because CMiPS is not rated for depths greater than 2000 m.  This gives two CMiPS profiles in deep water.


The vertical temperature and salinity profiles at the two stations were similar.  Both showed a well mixed layer just at the freezing point that extended from the surface to about 60 m to 80 m.  Below this layer, temperature and salinity increased as Upper Circumpolar Deep Water was encountered.  Both casts were deep enough to also include Lower Circumpolar Deep Water, which was seen at about 550 m to 600 m.


Water with temperature of 1.85ºC was encountered at 400 m at station 44 and 1.81ºC water was encountered at station 45 at 300 m.  These temperatures indicate that the southern boundary of the Antarctic Circumpolar Current was crossed.  This is consistent with observations from other outer shelf stations that suggest that the southern part of the Antarctic Circumpolar Current is flowing along the outer edge of the west Antarctic Peninsula continental shelf.


Sea Birds (Chris Ribic and Erik Chapman)

Surveys were conducted on 21 August for over 7 hours as the ship traveled between stations 44 and 45.  This transit was just off the continental shelf and ran roughly perpendicular to the shelf break.  Ice was mostly vast floe in 10/10ths ice concentration.  Temperatures were around 20ºC, and wind-chill dipped into the high 40's.  The transect crossed several large leads that had recently frozen over and were now covered by a thin layer of new gray ice.  The ship's track snaked through many large ice-bergs and second year sea-ice floes. 


During the previous evening, First Mate Mike Watson saw an Emperor Penguin at station 44.  An Emperor was also seen during the survey alongside a recently frozen lead.  A steady stream of Snow Petrels was observed flying directionally past the ship.  All of the birds were flying in the same general direction, moving from inshore toward the ice edge.  Perhaps these birds are moving from leads that have frozen over in the recent cold temperatures to feed in open water along the ice edge.  An adult Southern Giant Petrel was also observed flying in the direction of the ice edge.  No Crabeater Seals or Adélie Penguins were observed in the survey period, though penguin and seal tracks were seen on the surface of the ice.


A summary of the birds and marine mammals observed on 21 August (YD 233) during 7 hours, 14 minutes of survey time as the ship traveled between stations 44 and 45 is the following:


Species (common name)

Species (scientific name)

Number observed          

Snow Petrel            

Pagedroma nivea                


Emperor Penguin        

Aptenodytes forsteri            


Southern Giant Petrel  

Macronectes giganteus           




Marine Mammal report (Chico Viddi)

“Unpredictable, changing, wonderful, and wild”, these are the words that come to mind when thinking of Antarctica. The last three days we have seen more colors, felt the strength of Antarctic winds and felt more cold than all other days during this cruise. A sunny day on 19 August brought us winds of 35 to 45 kts, with gusts of 50 kts (at 1513).  It was also the coldest day since we left Punta Arenas, reaching a wind chill of -59.2ºC. On Tuesday the 20th, wind strength went down, but it was cloudy, dark, grey, and foggy. These two days had mainly ice floes of first year ice, varying from 8 to 10/10ths coverage, with some variable sized leads of open water or fresh ice. Few effort hours were made during 19 and 20 August, since we spent most of the day at stations and weather conditions and visibility were not optimal for a survey. Even though there were 15 hours of observation, only 2 hours corresponded to effective effort. During the entire period, no cetaceans were seen.  Only two crabeater seals were observed while at station 43 on 20 August. On August 21, this beautiful place gave us great presents, starting with a wonderful moonset, followed by a red-pink sunrise. Today's sighting conditions were excellent, even though wind speed reached 35 kts and air temperature and chill wind values were very low (-24 and 53ºC respectively). Three whale sightings were made. The first was by Captain Joe at 1054 (-67º 0.51′S; 73º 22.25′W) while headed to station 45 (identification is unclear). Later, two Minke whales were observed in two different sightings, at 1413 (-67º 02.88′S; 73º 33.51′W) and 1526 (-67º 08.49′S; 74º 11.34′W). For the first two sightings, the whales were swimming in a wide lead of open water (100-200 m); for the third, the whale was seen surfacing in a very narrow opening (less than 50 m).  The day could not have ended better, with a sunset and moonrise that colored clouds, icebergs, and pack ice with marvelous reds, oranges, and pinks.


Krill distribution, physiology, and predation (Kendra Daly, Kerri Scolardi, Emily Yam and Jason Zimmerman)

We completed two more Tucker trawls at stations 44 and 45.  The composition of zooplankton in the upper 100 m in a very diffuse scattering layer was similar at both sites.  The euphausiid, Thysanoessa macrura, and the copepod, Calanus propinquus, were relatively abundant.  In addition, we collected a few Euphausia superba, other copepods (e.g., Metridia, Paraeuchaeta), some salps, a few ctenophores and amphipods, and a polychaete worm.


ROV report (Scott Gallager, Phil Alatalo, Alec Scott)

ROV under-ice survey # 9 was conducted at station 43 on 20 August. This region was characterized by small, 1-m thick, re-frozen ice floes about 10 to 20 m in diameter with an extensive ridge field on the order of 1 to 2 m high. These characteristics proved difficult for the ROV to operate in because of the dangers of tether entanglement on the extensive 3-dimensional ice structure below. The bridge also had a difficult time finding stable floe on which the ice team could be deployed because cracks appeared quickly after the ship began to penetrate an area. After three attempts, a solid floe was located and we were off. First the ice team landed on the ice and then the ROV went into the water at 0727. Before the deployment, the bulb for the high intensity strobe on the 3D camera needed to be replaced. This required about 20 minutes of extra time on the deck allowing the temperature of the vehicle and cameras to reach the air temperature of 23ºC. Upon entering the water, ice crystals formed on the metal parts and roared to the surface in a turbulent plume. The glass camera windows iced over with an opaque layer which looked like someone had painted the windows black. The ROV was picked up out of the water, the camera windows sprayed with ethanol to dissolve the ice, and then returned to the water. The first hour of the dive was spectacular as the ROV ascended into cavernous under ice-domed structures constructed of rafted floes. As we moved through our transects, there were scattered furcilia up in the shallowest regions and small groups of 5 to 10 individuals located at 8 to 10 m depth immediately below the deepest ridges. Small numbers of ctenophores drifted by followed by the occasional fish larva. The under-ice surface of the floe being worked on by Frank Stewart and the ice team was smooth with the exception of pock marks 10 cm or so in diameter. A small number of single furcilia floated by with no signs of aggregations. The ridge line was clearly visible on the ROV mounted sonar system allowing good measurements of floe and ridged surfaces.


The ROV then moved into some unconsolidated brash and bergy bits on order 2 to 10 m in diameter. Furcilia were present in low numbers (approximately 2 to 10 individuals per cubic meter). The brash was moving around a bit with a 1 kt current.  Soon the ROV could not be moved up, down, or laterally within this minefield of ice. The tether could be seen to have lodged into several crevasses in front of the vehicle and additionally behind the vehicle in its blind spot, keeping it on a short leash. We worked for several hours maneuvering in and around the ice blocks and ridges trying to untangle the trapped tether. A clump weight was dropped down the tether from the ship in an attempt to lower the point of insertion and free at least one end. Several times during this ordeal we gained some tether giving us the false impression that we had been freed; however, as one region of the tether was freed, additional entrapment points became evident. We were faced with two tenable options: Send in the divers who might be able to free the tether, but possibly at the same time jeopardize their own safety in this difficult environment. Another thought, provided by Chief Scientist Peter Wiebe, was to very gently turn the ship screws and get the water moving behind the ship just enough to cause the brash to move and possibly free the tether without damaging the vehicle.  We decided to try the second idea.  Mate Mike Watson deftly engaged the screws and put on 5º pitch at low rpm. After a minute, we could see from the cameras on the ROV that the brash was moving. The deck crew continued to tug on the tether and the ROV dove hard. Suddenly, we could see daylight off in the distance and we knew we were free. The ROV was retrieved, but sustained some damage to a thruster due to being jammed with ice. A spare thruster was swapped for the damaged one and we were ready to go by the time we arrived at the next station.


ROV deployment 10 on 21 August at Station 44 was cut short because of the current overload indicator lights on two thruster motors.  The starboard and lateral motors showed that a water leak had probably occurred and needed immediate attention. The ROV was retrieved 20 minutes after entering the water, but not before making observations on furcilia abundance close to the ship. Approximately 1 to 2 individuals were present most of the time the vehicle was close enough to the surface to have the 3D cameras in focus.  No observations were made at depth because we were making preparations for immediate retrieval.  Following retrieval, the two offending thruster motors were removed, cleaned, and refilled with a special oil to keep water from entering the shaft housing.


BIOMAPER II group report (Gareth Lawson, Peter Wiebe, Scott Gallager, Phil Alatalo Dicky Allison, Alec Scott)

At 1330 on 21 August, we deployed the BIOMAPER-II for a four hour towyo through the deep waters between stations 44 and 45. The only acoustic feature present was a diffuse shallow scattering layer that was as much as 80 m thick and generally associated with the pycnocline. Near the start of the tow, this layer was centered at about 110 m, but by 1730 appeared to have moved to a more shallow distribution centered at 50 m, possibly indicative of a diurnal vertical migration. Images captured by the VPR suggest that this layer was composed of copepods, including large calanoids. Interestingly, we also captured images of numerous diatom chains, large single cell diatoms, and large radiolarians. The cameras on the VPR have an image volume of approximately 1.6 ml for the high magnification camera and 20 ml for the low magnification camera. The imaging rate is 60 images per second giving volumes sampled of 6 liters per minute and 72 liters per minute, respectively. As such, for us to catch an image of a planktonic organism they must be relatively abundant, typically greater than about 20 per cubic meter. In dense patches or aggregations of copepods or krill, the VPR captures images at a rate suggesting their nearest neighbor distance to be on the order of 1 cm or less. We have found such densities to be fairly rare this year compared with previous cruises in the fall and winter months.


In order to contribute substantially to our measurements of acoustic backscatter, organisms must either be quite large relative to the wavelength of the incident sound wave, or must occur at extremely high densities. As such, the small diatoms and copepods captured by the VPR may not be responsible for all of the backscatter we are observing at our currently operating frequencies of 120 and 200 kHz, unless they are present at very high densities. What might also be contributing to the remaining backscatter remains unclear, but we hope that the MOCNESS net samples will shed some light on this question.


In the past few BIOMAPER-II towyos, we have been plagued by problems with the tow-body not flying properly. In order to collect the best data possible, we like the body to remain perfectly level, but recently it has been flying at as much as a 22º angle nose-upward and 22º to port. Our initial attempts at solving this problem involved angling the adjustable aileron on the tail fin slightly downwards and adding horizontal stabilizers mid-way up the fin. Since these measures failed to solve the problem, most recently we have replaced the plastic tailfin used since the start of the cruise with a plywood one made last winter. The plastic fin had a strong curve to starboard, whereas the woods one is much straighter, so we are hoping this will solve the problem. Problems with electrical noise on the VPR video signals have been solved for the most part as we become poised to sample the offshore stations of 45, 46 and 64.


Current Position and Conditions

The broad-scale survey work for the past couple of days on the Palmer has been focusing on doing the work at the off-shore stations in the central sector of the Southern Ocean GLOBEC grid.  This is because we have been able to effectively work in the pack ice in this area.  We are currently steaming from station 46 to station 64 while towyoing BIOMAPER-II. Our position at 0017 on 22 August is -67º 30.098′S; -74º 20.048′W.  The air temperature is -26.7ºC and the barometric pressure is 989.7 mb. Winds are around 15 kts out of the southwest (230). Skies are clear and the pack ice is mostly 10/10.


Cheers, Peter